OpenStudy (anonymous):

Find exact coordinates of the centroid of the region bounded by the given curves. y = 4sin(x) y = 4cos(x) x = 0, x = π/4

OpenStudy (anonymous):

this one, im 100% sure i got the right answer, but it keeps saying im wrong :( no bueno

OpenStudy (bahrom7893):

My = ∫ [a to b] x [f(x) - g(x)] dx Mx = 1/2 ∫ [a to b] [f(x)^2 - g(x)^2] dx M = ∫ [a to b] [f(x) - g(x)] dx Are those the formulas?

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20100224155807AAjGmzy welll il be damned.

OpenStudy (bahrom7893):

HAHA!

OpenStudy (anonymous):

damn it all. -sigh-